OpenStudy (anonymous):

25xy^4 + 60xy^3/5xy= ?

zepdrix (zepdrix):

\[\huge 25xy^4+\frac{60xy^3}{5xy}\]Is this what the problem looks like? Only the second part is a fraction right?

OpenStudy (anonymous):

the 25xy^4+60xy^3 is supposed to be over 5xy

zepdrix (zepdrix):

Oh ok :)

OpenStudy (anonymous):

I've tried doing it over and over but none of the answers I get are one of the multiple choice answers.

zepdrix (zepdrix):

\[\large \frac{25xy^4+60xy^3}{5xy}\]We can write this as two separate fractions,\[\large \frac{25xy^4}{5xy}+\frac{60xy^3}{5xy}\]Let's divide the constants in front first,\[\large \frac{5xy^4}{xy}+\frac{12xy^3}{xy}\]That part make sense?

OpenStudy (anonymous):

Yeah,

zepdrix (zepdrix):

Both terms have an x on top and an x on bottom, let's divide those out,\[\large \frac{5y^4}{y}+\frac{12y^3}{y}\]Then we can divide the y terms, they have the same base, so we'll SUBTRACT the exponents. Think of the y's on the bottom as y^1.\[\large 5y^{4-1}+12y^{3-1}\]Which gives us an answer of,\[\huge \color{purple}{5y^3+12y^2}\]

zepdrix (zepdrix):

Is that one of your choices or no? :O

OpenStudy (anonymous):

Yes, oh my gosh. Thank you! I figured out where I got mixed up. I forgot to subtract the exponents haha.

zepdrix (zepdrix):

ah :) heh Wait wait wait... are you the girl from the Wendy's commercials? :O

OpenStudy (anonymous):

No.. Hahaha

zepdrix (zepdrix):

Sorry I couldn't resist :3 lol

OpenStudy (anonymous):

That was the first time I heard that one, at school people call me Ariel -.-

zepdrix (zepdrix):

Hahaha nice XD Unda Da seeeaaaaa!